|
|
2.1.1 Narrative
Oversight of the response was transitioned from OSC Stephen Ball to OSC Kevin Eichinger on Saturday, March 7, 2015.
March 8, 2015
Most personnel were given the day off to recover but a skeleton crew accomplished the following tasks in a short day. The property owner confirmed that there was approximately 128 fluid ounces of mercury stored in two separate containers in the house. We have found both of the storage containers, however they were both empty.
Basement Screening Levels and Removal Activities:
- Morning screening levels throughout the mother-in-law suite were between 20,000 and 25,000 ng/m3 at a temperature of 35 to 40 degrees and ventilation through three open windows. The negative air machines were operational throughout the night.
- Levels throughout the garage were ~20,000 ng/m3 prior to opening the garage in the morning.
- Mercury beads were observed in the closet under the stairs in the playroom and in the playroom against the back wall (near the corner where puddles were previously discovered in the wall).
- The playroom and Guest bedroom 2 were vacuumed today to remove remaining sulfur and tile dust.
- The wall between the playroom and guest bedroom 2 was identified as a hot spot near the floor and the concrete floor in guest bedroom 2 was also identified as a hot spot. Operations tomorrow will attempt to address these two hotspots through removal and sealing.
Upstairs Screening Levels and Removal Activities:
- Morning entry indicated mercury concentrations throughout the upstairs between 1,000 and 3,000ng/m3 with ventilation throughout the night and one fan pushing fresh air into the upstairs. The temperature in the house this morning was 42 degrees.
- Item removal began in the dining and living rooms today.
- Ambient air levels in the house after lunch were around 5,000 ng/m3. The floor is reading between 10,000 to 17,000 ng/m3 in many locations throughout the living room and dining room.
Item Decontamination Activities:
- The heat tent for upstairs item is under construction.
- 398 bags of items have been removed and inventoried from the house.
510 Cove Road:
- No 510 Cove Road operations occurred during this period.
March 9, 2015
Basement Screening Levels and Removal Activities:
- Morning screening levels throughout the mother-in-law suite were between 15,000 and 20,000 ng/m3 at a temperature of 45 to 50 degrees and ventilation through three open windows. The negative air machines were operational throughout the night.
- Levels throughout the garage were ~20,000 ng/m3 prior to opening the garage in the morning.
- The wall between the playroom and guest bedroom 2 was removed today due to hotspots located on the floor around the wall. Electrical circuits were in the wall and the breaker was turned off during critical points of the removal.
- The remaining area of the basement was vacuumed today. Following vacuuming we screened again through the basement and identified additional hotspots.
- The stairs leading from the basement to the upstairs was identified previously as having beads on them so we vacuumed the stairs today. Following the morning vacuuming there were more beads identified including in the cracks between the riser and the tread. A second, more thorough vacuuming was completed followed by spot checks and vacuuming with an eye dropper to get in cracks where possible.
- The toilet and vanity were removed in preparation for applying leveling compound to the floor.
Upstairs Screening Levels and Removal Activities:
- Morning entry indicated mercury concentrations throughout the upstairs between 4,000 and 7,000ng/m3 with ventilation throughout the night and one fan pushing fresh air into the upstairs. The temperature in the upstairs of the house this morning was 54 degrees.
- Item removal continued in the dining and living rooms today.
- During lunch rooms were closed off and measured following lunch. The kitchen was the highest room at 23,000 ng/m3 (~600) and the nursery and guest bedroom 1 were between 6,000-8,000 ng/m3.
- Beads were found on the upstairs floor near, but not on, the stairs.
Item Decontamination Activities:
- The heat tent for upstairs item is complete
- More than 450 bags of items have been removed and inventoried from the house.
- Many of the bagged items upstairs are screening lower than 50,000 ng/m3 and are being placed in the heat tent for decontamination. Items that screen greater than 50,000 ng/m3 are being disposed of unless they are deemed to be sentimental.
Community Involvement Activities:
- EPA CIC's canvassed the surrounding homes to pass out a fact sheet and inform neighbors of future road cleanup operations later this week.
March 10, 2015
OSC met with the property owners to discuss restoration plans for the lower level and replacement plans for personal effects. Received permission to dispose of large pieces of furniture and various items. Discussed other containers of mercury that were potentially removed from the home by the grandson. Requested that property owner speak with her daughter and attempted to accounted for/assure they are in the custody of an adult. Advised that we could disposed of this mercury.
Basement Removal Activities:
- All ceiling tiles in the lower level living space were removed.
- The lower living space were heated and vented to identify hot spots. Additional hot spots were identified in bedroom 3. The wall between bedroom 3 and the utility room was removed. The area was vacuumed to remove additional mercury. It is believed that all hot spots have been identified on the lower level.
- A small section of the lower level living room concrete floor that was screened at >50,000 ng/m3 after multiple cleanings and treatments was coated with a floor leveling compound. This area will be screened to assure that planned coating/sealing operation is effective.
- The garage was heated and vented to identify additional hot spots. No additional hot spots found.
- The floors and walls in the garage were prepped for the two part epoxy sealer. The garage walls were sealed with a two part epoxy paint.
Upstairs Screening Levels and Removal Activities:
- Item removal finished in the kitchen, dining and living rooms today. Additional free mercury was found in a laundry container. The washer and dry were removed. Item removal in the master bed room started.
- Crews continued vacuuming the living room and laundry room. There a visible mercury micro-beads in the joints between the boards of the hardwood floor. A test was done with mercury decontamination solution.
Item Decontamination Activities:
- Continued moving bagged items into the heat tents for decontamination
- Many of the bagged items upstairs are screening lower than 50,000 ng/m3 and are being placed in the heat tent for decontamination. Items that screen greater than 50,000 ng/m3 are being disposed of unless they are deemed to be sentimental.
- Had the property owner identify items in the lower heat tents that could be disposed of. These items are proving difficult to decontaminate.
Road and Driveway Removal Activities:
- OSC screened the road. Areas of the road that were less than 6,000 ng/m3 was considered clean. Used the mercury vapor analyzer and a 1’x1’ grid to clear portions of the road. Treated the road surface with a mercury decontamination solutions and applied a mercury amalgam powder to areas with visible beads and on areas that exceed 6,000 ng/m3. The decontamination solution and the amalgam was removed with a wet/dry HEPA vacuum that will be disposed of at the end of the project. This technique was effective on the road surface. Road operations only commence if there is no impact to interior operations. Due to heating and venting operations extra ERRS members were available to assist the OSC with this operation. Confirmation of decontamination effectiveness was determined by a transfer test with footwear. Area was cleared when the footwear screened at less then 6000 ng/m3 after walking the area.
Community Involvement Activities:
- OSC met with neighbors adjacent to the site to discuss the current plans and timeline. Surrounding community is very concerned about the situation.
510 Cove Road:
- Crews began to install the carpet. Carpet install should be complete tomorrow. Installers needed to glue the tack strips down and time is needed for the glue to dry.
March 11, 2015
Met with the property owner and continued discussions about potential Hg that was removed from the property. Property owner stated that she believes that there is possibility that Hg was removed. Property owner will discuss the situation with her daughter and grandson and attempt to find out where the Hg is located.
Basement Removal Activities:
- One additional section of wall was removed. The entire lower level living space was vacuumed one additional time.
- The sealing/coating test in the lower level living room was successful.
- At approximately 0900 hours, the heating and venting operations began in the lower level living space and garage. A temperature of around 100 degrees F was achieved in the areas. Heating/venting operations continue through the day and through the night. Two ERRS personnel are onsite after-hours to perform a fire watch, maintain the heaters and monitor carbon monoxide levels. They will turn off the heaters at approximately 0700 hours on 3/12/2014 to allow the areas to vent and cool down for testing when the crews arrive back on-site at 0800 hours on 3/12/2014.
- One coat of epoxy sealer was applied to the garage floor.
- Filters in the HVAC equipment were removed and the system is currently running during the heating/venting operation.
Upstairs Screening Levels and Removal Activities:
- Item removal continued in the master bedroom and master bathroom.
- The decontamination test proved successful on the living room hardwood floors. ERRS crews completed three washes with the mercury removal solution and a wet/dry vacuum with a squeegee attachment in the kitchen, dining room, living room and hallway. A fourth treatment was also completed using a mercury decontamination solution in all of these areas. At approximately 1700 hours, the heating and venting operations began in the upper level living space and garage. Two ERRS personnel are onsite after-hours to perform a fire watch, maintain the heaters and monitor carbon monoxide levels. They will turn off the heaters at approximately 0700 hours on 3/12/2014 to allow the areas to venting and cool down for testing when the crews arrive back on-site at 0800 hours on 3/12/2014.
- ERRS crews completed three washes with the mercury removal solutions and a wet/dry vacuum with a squeegee attachment on the exterior deck. A fourth treatment was also completed using a mercury decontamination solution in this area.
Item Decontamination Activities:
- Continued moving bagged items from the master bedroom and master bath room into the heat tents for decontamination. Many of the bagged items upstairs are screening lower than 50,000 ng/m3 and are being placed in the heat tent for decontamination. Items that screen greater than 50,000 ng/m3 are being disposed of unless they are deemed to be sentimental.
- Disposed of property identified by the property owner or items that exceeded 50,000 ng/m3.
- Approximately 560 bags have been generated.
- Lawn equipment and other miscellaneous outdoor tools where moved from the heat tent to an outdoor shed.
- Property was consolidated into one tent. Two of the lower level tents were dismantled. A third tent was re-purposed for response equipment storage. One additional tent was constructed on the upper level in order to accommodate person effects from the master bedroom and master bathroom.
510 Cove Road:
- Carpet installation was completed. Key will be returned to the rental company tomorrow, 3/12/2015. No addition response actions planned at this location.
2.1.3 Enforcement Activities, Identity of Potentially Responsible Parties (PRPs)
2.1.4 Progress Metrics
Waste Stream |
Medium |
Quantity |
Manifest # |
Treatment |
Disposal |
Non RCRA Solid Waste |
Solid |
5 - 30 yard dumpsters |
|
|
Landfill |
RCRA Hazardous Mercury Contaminated Debris |
Solid |
5 cubic yards |
|
|
Landfill |
Elemental Mercury |
Solid |
27 pounds |
|
Reclamation |
|
|